OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

m= 3-(-7)/-1-4=-2 m=-2

OpenStudy (anonymous):

I second that

OpenStudy (anonymous):

Slope = rise over run. Up 10 back 5 so 10/ -5. Which is -2.

OpenStudy (s3a):

rise/run = Δy/Δx = (y_final - y_initial)/(x_final - x_initial) = (-7-3)/(4 - -1) = (-10)/(5) = -2 1. The Δ means "change" in mathematics. 2. I chose the first point you gave as the initial point, and the second point you gave as the final point (but you could have chosen the points in the different order like AtefehAnsari did - it always gives the same final answer as long as you remain consistent about what is the initial point and what is the final point); you're going from one point to another. How much you need to go in each direction is known as the slope.

OpenStudy (anonymous):

So you changed your mind about me having a metal? :c